# AnyMaker: Zero-shot General Object Customization via Decoupled Dual-Level ID Injection

### v3

@@ Abstract
Text-to-image based object customization, aiming to generate images with the same identity (ID) as objects of interest in accordance with text prompts and reference images, has made significant progress. However, recent customizing research is dominated by specialized tasks, such as human customization or virtual try-on, leaving a gap in general object customization. To this end, we introduce AnyMaker, an innovative zero-shot object customization framework capable of generating general objects with high ID fidelity and flexible text editability. The efficacy of AnyMaker stems from its novel general ID extraction, dual-level ID injection, and ID-aware decoupling. Specifically, the general ID extraction module extracts sufficient ID information with an ensemble of self-supervised models to tackle the diverse customization tasks for general objects. Then, to provide the diffusion UNet with the extracted ID as much while not damaging the text editability in the generation process, we design a global-local dual-level ID injection module, in which the global-level semantic ID is injected into text descriptions while the local-level ID details are injected directly into the model through newly added cross-attention modules. In addition, we propose an ID-aware decoupling module to disentangle ID-related information from non-ID elements in the extracted representations for high-fidelity generation of both identity and text descriptions. To validate our approach and boost the research of general object customization, we create the first large-scale general ID dataset, Multi-Category ID-Consistent (MC-IDC) dataset, with 315k text-image samples and 10k categories. Experiments show that AnyMaker presents remarkable performance in general object customization and outperforms specialized methods in corresponding tasks. Code and dataset will be released soon in https://github.com/LingjieKong-fdu/AnyMaker.
Figure 1: Diverse applications of AnyMaker. Given text prompts and one single reference image, our AnyMaker can achieve various customization for general objects with high ID fidelity and flexible text editability, without further fine-tuning.
@@ 1 Introduction
Recent advancements in diffusion-based text-to-image generative models [1, 2, 3] have entitled common users to create high-fidelity photo-realistic images with minimal expertise. However, generating images from specific user elements, such as a personal dog of users, remains a considerable challenge [4, 5, 6, 7] due to the difficulty of preserving the ID of specific elements. To address this limitation, object customization, which leverages reference images and textual descriptions to generate images that shares the same identity as objects of interest, has become a crucial topic.
Current object customization methods fall into two main categories: object-specific methods and object-agnostic methods. Object-specific methods, such as DreamBooth [8] and DisenBooth [9], inject customized knowledge into pretrained diffusion models via learning additional reference-specific identifiers. While these methods can be adapted to various objects, their effectiveness relies on time-consuming fine-tuning on a few reference images, which leads to the low efficiency of such methods. Object-agnostic methods, or zero-shot methods like PhotoMaker [6] and InstantID [5], on the other hand, employ large-scale training datasets to generate customized images without further fine-tuning. However, these methods predominantly concentrate on specialized tasks that are only applicable to specific domains, such as human customization or virtual try-on. In this paper, we target at zero-shot general object customization which has seldom been studied in previous works.
To this end, we propose AnyMaker, an innovative zero-shot general object customization framework in text-to-image generation tasks, which can achieve high-fidelity customization for general objects given a text prompt and one single reference image without any test-time fine-tuning. The efficacy of AnyMaker comes from its novel general ID extraction, dual-level ID injection, and ID-aware decoupling.
Specifically, in contrast to prior approaches [6, 7, 5] that merely utilize a single model to extract object representations, we employ an ensemble of self-supervised models to extract more comprehensive information from general objects, ensuring the extracted ID-related information is sufficient enough for the high-ID-fidelity customization. Then, to integrate the extracted ID information into the diffusion UNet while preserving the text editability of the model, we design a global-local dual-level ID injection strategy. At the global level, we fuse the class tokens of the extracted feature with the class word (such as dog or cat) of the text description, injecting semantic-level ID information into the diffusion UNet without damaging the other elements in the text. At the local level, we inject the patch tokens of the extracted feature into the diffusion UNet through newly added cross-attention modules without compressing the feature dimension, ensuring that the model receives a substantial amount of ID-related details, which is crucial for achieving high-ID-fidelity generation. Furthermore, we find that the ID information is usually coupled with redundant non-ID elements such as the motions and orientations of objects. Such coupling phenomena can disrupt the text editing ability and ID retention ability of the model, and further hinder the diversity of generation. To address this issue, we design a novel ID-aware decoupling module at the training stage, compelling the model to discern the intrinsic properties of the ID information and exclude non-ID information away from the generation process.
Benefiting from the innovative ID processing modules mentioned above, AnyMaker realizes zero-shot customization for general objects with high ID fidelity and flexible text editability. We demonstrate various applications of AnyMaker in Fig.1, such as realistic human customization, cartoon character customization, virtual try-on, story generation, ID mixing, and so on.
To validate our method and facilitate the development of general object customization, we build the first large-scale general ID dataset, that is Multi-Category ID-consistent (MC-IDC) dataset, which contains 315k text-image samples with over 10k categories. Experiments demonstrate that AnyMaker achieves state-of-the-art generation results on general object customization and even outperforms task-specific methods in specialized domains like human customization and virtual try-on.
In summary, the contributions of our work are as follows:
  • We propose a novel zero-shot general object customization framework in text-to-image generation tasks, which is capable of customizing general objects with high ID fidelity and flexible text editability through innovative general ID extraction, dual-level ID injection and ID-aware decoupling.
  • We construct a large-scale Multi-Category ID-Consistent (MC-IDC) dataset to significantly advance the development of general object customization.
  • We validate the effectiveness of our method through extensive experiments, which demonstrate that our AnyMaker delivers desirable performance in general object customization and surpass specialized methods tailored for specific tasks.
@@ 2 Related Work
Diffusion models. Diffusion models have demonstrated their effectiveness for text-to-image generations. The DDPM model pioneered in this field, using a diffusion and denoising process to create a mapping between Gaussian and image distributions. The Latent Diffusion Model (LDM) [10] took this a step further by applying the diffusion model to a latent space rather than pixel space, leading to the creation of text-to-image diffusion models like Stable Diffusion (SD), Midjourney, and DALLE-3 [11]. One main research problem for the diffusion model is its structure. Other than the original UNet, DiT  [12] and Pixart-$\alpha$ [3] adopt transformer as the backbone structure. Other researchers have also focused on enhancing image fidelity regarding complicated concepts. By manipulating the cross-attention maps according to text prompts,  [13] ensures that required objects are sufficiently generated in the images.  [14] proposed to process the text prompts to decompose all target attributes, which can be further used to guide the denoising procedure with the help of different objective functions. To enhance content consistency,  [15] adopted an inner localization method by calculating the cross attention. Our method is generally built on diffusion models. While different from these basic models for plain image generations, we focus on the more challenging topic of zero-shot any object customization.
Image customization. Image customization aims to generate images with the same objects of interest as the reference image [8, 16, 17, 18, 5]. Current methods can be generally grouped into two categories. One is object-specific model, which tries to finetune the pretrained diffusion models with reference images. Textual inversion [19] learns an extra word embedding for the target concept. Dreambooth [8] and DisenBooth [9] enhance such a pipeline with auxiliary objectives such as prior loss, which ensures the retention of the model’s original generative capabilities throughout the fine-tuning process. The other genre is object-agnostic models, which can directly process reference images without further fine-tuning. PhotoMaker [6], as a typical zero-shot method, is driven by training a person ID-extaction network, while  [20] uses a face model with CLIP to extract ID features. With the help of ControlNet [16], InstantID [5] can control face expression by leveraging a pre-trained face model to extract face embedding. MagicCloth [21] utilizes a ControlNet style garment extractor to encode garment details and fuse with attention for clothes replacement or preview. On the contrary to these works, we try to combine the versatility of object-specific models and the efficiency of object-agnostic models to present a method that can customize any object in a zero-shot manner, which has never been studied before.
@@
Figure 2: Overview of our AnyMaker. AnyMaker is a zero-shot text-to-image customization method for general objects, consisting of general ID extractor, global-local dual-level ID injection, and ID-aware decoupling module.
@@ 3 Method
We first outline the overall structure of AnyMaker in Sec. 3.1. Then, we detail the general ID extraction, dual-level ID injection and ID-aware decoupling in Sec. 3.2, Sec. 3.3 and Sec. 3.4 respectively.
3.1 Overview
Give a reference image and a text prompt, our goal is to generate images that shares the same ID as the object of interest from the reference image, and modifies non-ID elements such as motions and backgrounds in accordance with the text prompt. To this end, we propose AnyMaker, an innovative zero-shot text-to-image customization framework for general objects, as summarized in Fig. 2. Firstly, we utilize the general ID extractor to acquire sufficient ID information from the reference image with a segmentation mask of the interested object. Next, we leverage the dual-level ID injection to endow the diffusion model with ID information at the global and local levels respectively, without damaging the text editing ability of the model. Additionally, we design a novel ID-aware decoupling module to distinguish the ID information from the coupled non-ID elements in the object representation, for higher ID fidelity and better text editing ability.
3.2 Genral ID Extraction
We utilize an ensemble of self-supervised models, consisting of DINOv2 [22] and MAE [23], to extract the representations of general objects in the reference image.
Previous works [6, 5] targeting at customization tasks in specific domains like human face, usually use a single model to extract object representations. However, our experiments demonstrate that one single pre-trained model, such as CLIP [24], DINOv2 [22] or MAE [23], cannot extract sufficient features for high-ID-fidelity generation of general objects, as shown in Fig. 4 and Tab. 2. Specifically, previous works [4, 5] demonstrate that features extracted by CLIP contain insufficient details, thus being less practical as an ID feature extractor, while DINOv2 is capable of providing desirable details due to its self-supervised contrastive learning strategy. However, we find that the features extracted by DINOv2 are color-insensitive because it is trained with augmentations such as ColorJitter. Consequently, solely using DINOv2 as an ID extractor would lead to mistakenly generated color. MAE, on the other hand, is aware of the color thanks to its reconstruction-related objective.
Therefore, we leverage DINOv2 and MAE, which compensate for each other, to extract sufficient and comprehensive information to handle the challenging customization task for general objects with various categories and diverse characteristics. To align the extracted features and facilitate the subsequent injection, we adopt a two-layer MLP. The process of extracting general object representations is as follows,
@@ 3.2 Genral ID Extraction
where $\odot$ represents element-wise multiplication, and $f_{dino}^{C},f_{dino}^{P},f_{mae}^{C},f_{mae}^{P}$ represent the class tokens and patch tokens extracted by DINOv2 and MAE respectively. $F_{dino}$ and $F_{mae}$ represent the DINOv2 and MAE backbone. $I_{ref}$ represents the reference image, and $M_{ref}$ is the segmentation mask of the interested object in the reference image. Consequently, tokens derived from our general ID extractor not only offer intricate details from DINOv2, but also have sufficient colour and structural information from MAE.
@@ 3.3 Dual-level ID Injection
We propose a dual-level ID injection mechanism to inject the global semantic ID information and local fine-grained ID information into the diffusion UNet respectively, ensuring both flexible text editability and high ID fidelity in the customization process for general objects.
Global ID injection. In text-to-image generation, text prompts are leveraged to guide diffusion models to generate diverse images. Nevertheless, diffusion models are incapable of accurately generating images with a specific object ID solely relying on the text description. To inject the ID information of the interested objects into the text, while simultaneously maintaining the editing capability of text prompts for other elements such as the background and motions, we design the global ID injection mechanism. Specifically, we merge the semantic-level class tokens $(f_{dino}^{C},f_{mae}^{C})$ with the class word in the text embedding, such as "dog" or "cat",
@@
where $f_{fuse}^{C}$ is the fused class token, and $f_{text}^{C}$ represents the class word embedding of the interested object. Then we insert the fused class token $f_{fuse}^{C}$ to the original position of the class word in the text embeddings, obtaining the global-level condition $c_{g}$ with ID-related information of the object of interest, as shown in Fig. 2. Following this, the global-level condition $c_{g}$ engages with the cross-attention module in the UNet, like to standard text-to-image models [10], as follows,
@@
Local ID injection. Semantic-level information in the global-level condition $c_{g}$ contains relatively fewer details of the objects, and thus the model cannot handle complex customization tasks for general objects effectively merely relying on global ID injection. To supply the model with sufficient ID-related details, we propose the local ID injection mechanism. Specifically, we first fuse the extracted patch tokens $f_{dino}^{P}$ and $f_{mae}^{P}$ with a two-layer MLP to densify the contained information to a unified representation $c_{l}$, as follows,
@@
where $Q_{l}=Z_{l}W_{q}$, $K_{l}=c_{l}W_{k}$, $V_{l}=c_{l}W_{v}$ are the query, key, and value of cross-attention module, in which $W_{l}$, $W_{l}$, $W_{l}$ are the weight matrices of the trainable linear projection layers, and $Z_{l}$ is the query feature containing the latent input information.
Through the dual-level ID injection, we provide the model with sufficient ID-related information, both at the semantic level and the detail level, on the premise of not impairing the text editing ability of the model.

### v4

# CustAny: Customizing Anything from A Single Example

@@ Abstract
Recent advances in diffusion-based text-to-image models have simplified creating high-fidelity images, but preserving the identity (ID) of specific elements, like a personal dog, is still challenging. Object customization, using reference images and textual descriptions, is key to addressing this issue. Current object customization methods are either object-specific, requiring extensive fine-tuning, or object-agnostic, offering zero-shot customization but limited to specialized domains. The primary issue of promoting zero-shot object customization from specific domains to the general domain is to establish a large-scale general ID dataset for model pre-training, which is time-consuming and labor-intensive. In this paper, we propose a novel pipeline to construct a large dataset of general objects and build the Multi-Category ID-Consistent (MC-IDC) dataset, featuring 315k text-image samples across 10k categories. With the help of MC-IDC, we introduce Customizing Anything (CustAny), a zero-shot framework that maintains ID fidelity and supports flexible text editing for general objects. CustAny features three key components: a general ID extraction module, a dual-level ID injection module, and an ID-aware decoupling module, allowing it to customize any object from a single reference image and text prompt. Experiments demonstrate that CustAny outperforms existing methods in both general object customization and specialized domains like human customization and virtual try-on. Our contributions include a large-scale dataset, the CustAny framework and novel ID processing to advance this field. Code and dataset will be released soon in https://github.com/LingjieKong-fdu/CustAny.
@@ 1 Introduction
Recent advancements in diffusion-based text-to-image generative models (e.g., SDXL, Improving GAN, PixArt Alpha) [21, 3, 6] have empowered users to generate highly realistic images with minimal technical knowledge. However, generating images of specific objects [8, 30, 14, 33], such as personal pets, remains challenging due to the difficulty of preserving their unique identity. Object customization, a technique that utilizes reference images and textual descriptions to generate images of specific objects, has emerged as a critical area of research. Existing methods [27, 35, 20, 18, 30] for object customization, such as DreamBooth, PortraitBooth and InstantID, have limitations in general-purpose object customization: Object-specific models [27, 5], while powerful, are often inefficient due to the need for extensive fine-tuning for each new object. Conversely, object-agnostic models [14, 30] have limited capabilities and often struggle with diverse object types.
Unlike existing customizing applications tailored to one specific task like face generation or virtual try-ons, this paper focuses on general object customization from a single example, as shown in Fig. LABEL:fig:intro. Specifically, it addresses zero-shot object customization in various scenarios where no fine-tuning on specific objects is available. Achieving this first requires a large-scale dataset containing various categories of objects for model pre-training. At present, however, there are few publicly released datasets in the field of zero-shot customization for general objects. To address this issue, we propose a novel construction pipeline for general ID dataset and introduce the Multi-Category ID-Consistent (MC-IDC) dataset, the first large-scale dataset for zero-shot object customization across diverse scenarios. It includes over 315K high-quality images across 10K categories, such as human faces, animals, clothing, and tools, supporting a wide range of object types and domains. The dataset features reference-target image pairs with consistent object IDs, including segmentation masks and text captions, making it ideal for ID-consistent generation tasks and advancing research in general object customization.
Based on our MC-IDC, we design a zero-shot general object customization framework which is able to Customize Anything (CustAny) from a single example. The innovation of our framework is mainly reflected in managing and utilizing identity information beyond the narrow focus of previous work. Specifically, the CustAny framework follows three essential steps: ID extraction, ID injection, and ID disentanglement. These steps address three critical questions: (1) How to capture enough identity details to handle a variety of complex customization tasks? (2) How to add this identity information to the model while still enabling text-based edits? (3) How to reduce the impact of irrelevant features from the reference object on the final results?
Formally, CustAny features an ID extraction module, dual-level ID injection, and an ID-aware decoupling module, offering innovative ID handling methods. Unlike prior methods [14, 33, 30] that use a single model for object representations, we leverage multiple pre-trained self-supervised models to capture more detailed ID-related features for high-ID-fidelity customization. To integrate this information into the diffusion model, we use a dual-level ID injection strategy: globally, by merging semantic-level ID features with text descriptions, and locally, by injecting patch tokens through cross-attention modules, preserving both ID fidelity and text editability.
Furthermore, we identify that ID information is often mixed with redundant non-ID elements, like object motion and orientation, which can disrupt both text editing and ID retention. To solve this, we introduce an ID-aware decoupling module during training to help the model separate ID-related properties from non-ID elements, enhancing both ID fidelity and generation diversity. With these advanced ID processing modules, CustAny enables zero-shot customization for various objects while supporting flexible text-based editing. As shown in Fig. LABEL:fig:intro, CustAny applies to object customization in various scenarios such as human customization, cartoon character personalization, virtual try-on. We conduct extensive experiments showing that CustAny achieves state-of-the-art performance in general object customization and outperforms task-specific methods in areas like human customization and virtual try-on.
In summary, we have these contributions: (1) Dataset: We propose a novel construction pipeline for general ID dataset, and present a large-scale dataset, MC-IDC, with 315K images across 10K categories for zero-shot object customization. (2) CustAny Framework: We design a zero-shot framework for general object customization in text-to-image generation, focusing on ID fidelity and text-based editing. (3) ID Processing: CustAny uses multiple self-supervised models and a dual-level ID injection strategy to integrate ID information while preserving text-editability. We present a module to separate ID from non-ID elements, improving ID fidelity and generation diversity. (4) Good Results: CustAny outperforms task-specific methods in general and specialized object customization tasks, including human customization and virtual try-on.
Figure 1: Illustration of our general ID dataset, MC-IDC. In each sample, the reference image with the object mask provides ID information, the text prompt offers semantic-level guidance for generation, and the target image serves as the ground truth.
@@ 2 Related Work
Diffusion models. Diffusion models have demonstrated their effectiveness for text-to-image generations. The DDPM model pioneered in this field, using a diffusion and denoising process to create a mapping between Gaussian and image distributions. The Latent Diffusion Model (LDM) [25] took this a step further by applying the diffusion model to a latent space rather than pixel space, leading to the creation of text-to-image diffusion models like Stable Diffusion (SD), Midjourney, and DALLE-3 [24]. One main research problem for the diffusion model is its structure. Other than the original UNet, DiT [17] and Pixart-$\alpha$ [6] adopt transformer as the backbone structure. Other researchers have also focused on enhancing image fidelity regarding complicated concepts. By manipulating the cross-attention maps according to text prompts,  [4] ensures that required objects are sufficiently generated in the images.  [9] proposed to process the text prompts to decompose all target attributes, which can be further used to guide the denoising procedure with the help of different objective functions. To enhance content consistency, Tewel et al. [29] adopted an inner localization method by calculating the cross attention. Our method is generally built on diffusion models. While different from these basic models for plain image generations, we focus on the more challenging topic of zero-shot any object customization.
Image customization. Object customization techniques generally fall into two categories: object-specific and object-agnostic (zero-shot) methods. Object-specific approaches, like DreamBooth [27] and DisenBooth [5], fine-tune pretrained diffusion models by incorporating new, object-specific identifiers from a small set of reference images. While effective, these methods require extensive fine-tuning for each new object, which limits their scalability and efficiency. In contrast, object-agnostic or zero-shot approaches such as PhotoMaker [14] and InstantID [30], leverage large-scale training datasets to customize images without the need for additional fine-tuning. Although more efficient, these methods are often domain-specific, focusing on tasks like human customization, facial manipulation, or virtual try-on applications. For example, PhotoMaker trains a person ID extraction network, while InstantID applies ControlNet [35] to control facial expressions based on pretrained embeddings, and MagicCloth [7] uses ControlNet to customize garment details for virtual try-on. These existing methods, however, have limitations when it comes to general-purpose object customization. Object-specific models are powerful but inefficient, requiring time-intensive training for each new item, while object-agnostic models tend to be narrow in focus and are often limited to specific types of objects or applications. In this work, we aim to bridge these gaps by combining the versatility of object-specific methods with the efficiency of zero-shot methods. As illustrated in Fig. LABEL:fig:intro, our approach targets the ability to customize any object based on a single example image and text description, achieving generalization across diverse categories with minimal fine-tuning. This design marks a significant step toward general object customization across various domains, an area that remains largely unexplored.
@@
Figure 2: Overview of our CustAny. CustAny is a zero-shot text-to-image customization method for general objects, consisting of general ID extractor, global-local dual-level ID injection, and ID-aware decoupling module.
3 MC-IDC Dataset
We curate the Multi-Category ID-Consistent (MC-IDC) dataset, the first large-scale dataset for general object customization research. We present the overall illustration of the MC-IDC as shown in Fig. 1. The dataset’s merits and construction process are described below, with additional details in Supplementary.
Construction pipeline. The MC-IDC is constructed by the following four steps: (1) Data collection. MC-IDC dataset contains diverse sources including web-crawled images, movies and publicly available datasets. To facilitate the subsequent generation of image pairs, most of the data we collect are video datas or multi-view datas. To ensure the high quality of our dataset, we delete the images with a resolution less than 300$\times$300. (2) Instance detection and segmentation. For some publicly available datasets or web-crawled images which do not provide segmentation annotations, we use advanced instance segmentation model [31] for instance detection and segmentation. For movie dataset, we extract frames with 10fps and do object tracking with [19], aiming to establish the ID connection across frames. (3) Image pair generation. For most data samples which are from a clip of video or a group of multi-view data, we randomly select two frames containing the same object and perform a random crop around the object to form the reference-target image pair. For the other data samples from the single image dataset, we perform random data augmentations on the single image two times, and take them as the reference image and the target image respectively. (4) Text prompt generation. In each image pair, the target image is densely captioned with the state-of-the-art large vision language model [1], serving as the text prompt to guide the generation.
Merits of MC-IDC: Our MC-IDC dataset contains 315k high-quality samples across 10k+ categories and diverse domains, with each sample featuring reference-target image pairs, segmentation masks, and text captions to support ID-preserving, diverse object customization. Particularly, it has these merits: (1)Diverse domains and categories: Our MC-IDC dataset contains around 315k samples across 10k+ categories, including human faces, animals, clothes, and tools. It features images from diverse domains, such as real-world photos, animations, model-generated content, and movies, enabling general ID understanding. Further details are provided in the Supplementary. (2)High quality: With an average resolution of 1039$\times$950, our dataset ensures high-quality training and inference for our model. (3) Reference-target image pair: Each sample includes a reference image with a segmentation mask, a target image, and a text caption. The reference image provides ID information, while the target image and text guide semantic generation. The reference and target images depict the same object in different states, sharing the same ID and varying non-ID elements like motion and direction. This setup ensures ID preservation while maintaining generation diversity.
@@ 4 Method
Overview. Given a reference image and text prompt, our goal is to generate an image that retains the identity (ID) of the object in the reference while modifying non-ID elements, like motions and backgrounds, based on the prompt. Thus we introduce CustAny, a zero-shot text-to-image customization framework for general objects, as in Fig. 2. We first use an ID extractor to capture ID information from the reference image with a segmentation mask. Dual-level ID injection is then applied to embed the ID into the diffusion model globally and locally, preserving text-editing capabilities. Finally, an ID-aware decoupling module separates ID details from non-ID elements, improving ID fidelity and text-editing accuracy.
4.1 General ID Extraction
To extract comprehensive ID information, we use a combination of self-supervised models, DINOv2 [16] and MAE [12], to extract representations for general objects in the reference image. Specifically, previous methods, such as [14, 30], which target specific domains like faces, typically use a single model like CLIP [22] to extract object features. However, we find that using only one pre-trained model—whether CLIP, DINOv2, or MAE—leaves gaps in capturing features necessary for high-ID-fidelity in general objects (see Fig. 5 and Tab. 2). Essentially, while CLIP as in [8, 30] lacks sufficient detail, DINOv2 (despite its strong details due to contrastive learning) is color-insensitive because it is trained with color-augmented techniques like ColorJitter. MAE, however, retains color sensitivity due to its reconstruction-based training. Thus, combining DINOv2 and MAE allows us to capture both detail and color for more accurate ID extraction.
We combine DINOv2 and MAE, which complement each other, to capture comprehensive information for the challenging task of customizing diverse general objects. To align and prepare these features for injection, we use a two-layer MLP. The extraction process for general object representations is as follows,
@@ 4.1 General ID Extraction
where $\odot$ represents element-wise multiplication; and we have the class tokens and patch tokens $f_{dino}^{C},f_{dino}^{P}$, $f_{mae}^{C},f_{mae}^{P}$ extracted by DINOv2 and MAE respectively. $F_{dino}$ and $F_{mae}$ are the DINOv2 and MAE backbone. $I_{ref}$ denotes the reference image; and $M_{ref}$ is the segmentation mask of the interested object in the reference image. Consequently, tokens derived from our general ID extractor not only offer intricate details from DINOv2, but also have sufficient colour and structural information from MAE.
@@ 4.2 Dual-level ID Injection
We aim to inject maximum extracted ID information into the diffusion UNet without compromising its text-editing ability. Our dual-level ID injection mechanism separately embeds global semantic ID and local fine-grained ID into the UNet, ensuring flexible text edits and high ID fidelity in customizing general objects.
Global ID injection. In text-to-image generation, text prompts guide diffusion models to create diverse images. However, diffusion models cannot reliably generate images with a specific object ID based on text alone. To embed object ID information into the text while preserving editability for other elements like background and motion, we introduce a global ID injection mechanism. Specifically, we merge the semantic-level class tokens $(f_{dino}^{C},f_{mae}^{C})$ with the class word in the text embedding, such as “dog” or “cat”,
@@
where $f_{fuse}^{C}$ is fused class token and $f_{text}^{C}$ is class word embedding. We insert $f_{fuse}^{C}$ into the class word position in the text embeddings, obtaining the global-level condition $c_{g}$ with ID-related information, as in Fig. 2. The global-level condition $c_{g}$ interacts with cross-attention module of UNet, similar to standard text-to-image models [25] as,
@@
Local ID injection. The global-level condition $c_{g}$ contains limited object details, making it insufficient for complex customization tasks. To provide more ID-related information, we propose the local ID injection mechanism, where we fuse the extracted patch tokens $f_{dino}^{P}$ and $f_{mae}^{P}$ using a two-layer MLP to create a unified representation $c_{l}$,
@@
where $Q_{l}=Z_{l}W_{q}$, $K_{l}=c_{l}W_{k}$, $V_{l}=c_{l}W_{v}$ are the query, key, and value of cross-attention module, in which $W_{l}$, $W_{l}$, $W_{l}$ are the weight matrices of the trainable linear projection layers, and $Z_{l}$ is query feature containing the latent input information. By dual-level ID injection, we have the model with sufficient ID-related information at both semantic and detail levels, without impairing its text-editing ability.